# AudioTrust: Benchmarking the Multifaceted Trustworthiness of Audio Large Language Models

### v3

@@ 1 Introduction
Despite rapid progress in ALLMs (Du et al., 2023; Tang et al., 2024; Chu et al., 2024; Zhang et al., 2023a; 2024a; Xu et al., 2025c; Yao et al., 2024), there is still no comprehensive benchmark evaluating their potential risks. Existing safety evaluation frameworks, such as SafeDialBench (Cao et al., 2025) and SafetyBench (Zhang et al., 2023c), mainly focus on the text modality while overlooking the unique properties or application scenarios of audio. Unlike text, audio features complex temporal-frequency patterns, rich emotions, and contextual dependencies, and introduces additional challenges, such as speech privacy, speaker recognition, and environmental acoustic analysis. These factors lead to unique trustworthiness risks in audio, including gender and accent biases, audio hallucinations, social engineering attacks, personal information leakage, and adversarial attacks on audio systems.
The integration of audio modalities into large models, while functionally powerful, introduces a new attack surface and exacerbates existing trustworthiness vulnerabilities. To systematically quantify these emergent risks, we introduce AudioTrust, the first comprehensive benchmark designed to evaluate the trustworthiness of ALLMs. AudioTrust establishes a rigorous evaluation framework across six critical dimensions where audio introduces unique safety concerns: (1) Fairness: Evaluating biases derived from vocal delivery, rather than the semantic content of speech; (2) Hallucinations: Testing for audio-grounded hallucinations, where model outputs violate the physical laws or temporal logic of an acoustic scene. (3) Safety: Assessing resilience to harmful queries that leverage persuasive or emotional vocal tones to bypass safety filters; (4) Privacy: Quantifying the leakage of sensitive information from spoken content and inference of personal attributes from acoustic cues; (5) Robustness: Assessing model performance in acoustically complex and imperfect environments, such as those with background noise. (6) Authentication: Evaluating the ability to distinguish authentic speakers from sophisticated impersonation attacks, including voice clones and audio-based social engineering. Underpinning our benchmark is a curated dataset of over 4,420 audio samples spanning 18 distinct evaluation tasks, from emergency communications to adversarial settings. We deploy a large-scale automated evaluation pipeline to ensure rigorous and reproducible assessment. The reliability of our automated metrics and results is verified by human experts (over 97% agreement rate). The initial findings on representative models are summarized in a public leaderboard (see Figure˜2(b)). Details of the benchmark are provided in Appendix˜C and Figure˜4.
@@
Safety: Incorporating audio inputs substantially broadens the attack surface. Unlike text, speech carries emotional, contextual, and anthropomorphic cues that adversaries can exploit by modulating tone, injecting affective signals, or impersonating identities. We design a composite attack framework spanning emotion-driven deception and identity-verification evasion. Empirical analysis shows: (1) closed-source models exhibit stronger overall robustness but remain sensitive to highly emotional speech; (2) open-source models are disproportionately vulnerable to identity- and emotion-based manipulation, often yielding unsafe outputs in high-stakes settings such as healthcare; (3) for prompts concerning illicit guidance, closed-source systems largely resist, whereas some open-source models deliver risky recommendations in multi-turn audio dialogues. Further details are provided in Section˜3.
@@ 2 AudioTrust: Fairness
Attack Strategies. To systematically probe these risks, we categorize fairness into two dimensions: traditional fairness and audio-based fairness. Traditional fairness assesses biases linked to demographic attributes like gender, race, and age (Cho et al., 2023; Zhang et al., 2017b; Wan et al., 2024) that can be inferred from a voice. We test, for example, if a model’s loan approval decisions are skewed by whether the applicant sounds male or female. Audio-based fairness isolates biases triggered by instrinsic acoustic properties. We divide these biases into four sub-categories: accent, linguistic characteristic (e.g., speech disfluencies may adversely impact the model’s fairness), economic status (e.g., noisy environments, often correlated with lower economic status, might yield more negative outputs), and personality traits (e.g., negative emotions may cause the model to produce negative outputs) (see Figure 5). For each sensitive attribute, we designed decision-making and stereotype-driven scenarios (Xu et al., 2025a; Wang et al., 2024b). Decision-making covers recruitment, admissions, and loan evaluations. Stereotype-driven contexts address beliefs such as men outperforming women in mathematics and gender roles in medicine and occupations. We constructed a dedicated dataset consisting of 840 audio samples, each lasting approximately 20 seconds. The samples were annotated with seven key sensitive attributes: gender, age, and race for evaluating traditional fairness, as well as accent, linguistic characteristic, economic status, and personality traits for assessing audio-baseed fairness. Notably, due to the limitations of the audio modality, attributes such as appearance are not included (Nangia et al., 2020; Parrish et al., 2021). Detailed dataset construction procedures can be found in Sections˜D.1 and D.2.
Evaluation & Metrics. For fairness evaluation, we introduced a key metric: Group Unfairness Score (Xu et al., 2025a; Deldjoo & Nazary, 2024). Group unfairness examines the distributional equity of model outputs across different groups (e.g., male versus female), measured by the divergence or balance between group distributions. indicates perfect fairness, while indicates extreme unfairness. In computing the unfairness score, we used human annotation and counting throughout to ensure data correctness and validity. For detailed formulas in the Section˜D.3.
@@
| SALMONN | 0.861 | 0.911 |
| Ultravox | 0.762 | 0.608 |
| Qwen2-Audio | 0.667 | 0.710 |
| MiniCPM-o 2.6 | 0.740 | 0.585 |
| Step-Fun | 0.342 | 0.495 |
| Qwen2.5-Omni | 0.933 | 0.798 |
| Kimi-Audio | 0.964 | 0.914 |
| OpenS2S | 0.983 | 0.843 |
| Step-Audio2 | 0.926 | 0.902 |
@@
| Gemini-1.5 Pro | 0.703 | 0.540 |
| GPT-4o Audio | 0.074 | 0.736 |
| GPT-4o mini Audio | 0.136 | 0.755 |
| Gemini-2.5 Flash | 0.630 | 0.754 |
| Gemini-2.5 Pro | 0.681 | 0.795 |
| Average | 0.672 | 0.739 |
Note: : higher than column average, : lower than column average, subscript is signed difference from mean (3 decimals).
@@ 3 AudioTrust: Hallucination
Attack Strategies. In AudioTrust, we identify two main categories of audio hallucinations (see Appendix E.1): Physical Logic and Chronological Order. The former relates to violations of acoustic laws and environment properties, and the latter reflects failures in reasoning about temporal and causal relations. These distinctions directly connect to safety risks in real-world use. To evaluate them, we built a dataset of 320 samples from synthetic and real sources (see Appendix E.2). For Acoustic-Physical Hallucinations, we focus on two specific manifestations: (1) Contravention of Physicochemical Constraints, generating impossible events (e.g., the flames are burning in the seawater) to test propagation understanding (Liu et al., 2024); (2) Source–Environment Mismatch, applying contradictory reverberation (e.g., casual speech with cathedral acoustics) to test disentanglement of source vs. environment (Zhou et al., 2022). For Temporo-Logical Hallucinations, we examine (1) Temporal-Causal Inversion: reversing causal chains (e.g., engine start before ignition) to probe event logic (Chen et al., 2023); (2) Cross-Modal Contradiction: pairing audio with conflicting text (e.g., fast footsteps described as peaceful rain) to test cross-modal reasoning (Deshmukh et al., 2023).
Evaluation & Metrics. We introduce a comprehensive evaluation framework to assess model capabilities across four key dimensions: hallucination detection, attribute verification, real-world consistency, and transcription accuracy. For hallucination detection, models are required to identify inaccuracies in audio-text pairs and provide justifications. Performance is quantified via a multi-dimensional, GPT-4o-based evaluator (OpenAI et al., 2024).The scores mentioned above range from 0 to 10 and are used to evaluate the accuracy of the detection and the quality of the interpretation. These scores were then subject to further review by the human evaluators. We probe for attribute-level hallucinations related to physical properties, labels, and content using multi-stage prompting (Madaan et al., 2023). To measure alignment with factual information, we adopt the two-stage protocol from Li et al. (2023a) for real-world consistency assessment. Finally, we evaluate transcription robustness under hallucinatory interference using both the standard Word Error Rate (WER) (Deshmukh et al., 2023) and the cross-modal WER (CM-WER) (Tang et al., 2024). The complete experimental design and metric details are provided in Appendices E.3.1 and E.3.2.
@@
| MiniCPM-o 2.6 | 6.24 | 6.20 | 8.28 | 6.13 |
| Qwen2-Audio | 8.15 | 4.34 | 7.26 | 7.77 |
| SALMONN | 2.65 | 1.22 | 6.64 | 3.98 |
| Ultravox | 5.74 | 4.52 | 8.01 | 8.34 |
| Qwen2.5-Omni | 8.12 | 5.63 | 7.89 | 6.11 |
| Step-Fun | 3.96 | 4.84 | 5.80 | 8.72 |
| Kimi Audio | 1.86 | 5.77 | 5.82 | 8.54 |
| Step-Audio2 | 3.62 | 1.94 | 2.60 | 2.76 |
| OpenS2S | 1.92 | 5.03 | 5.97 | 7.89 |
@@
| Gemini-1.5 Pro | 8.41 | 7.81 | 8.66 | 8.87 |
| Gemini-2.5 Flash | 7.98 | 8.36 | 8.71 | 8.57 |
| Gemini-2.5 Pro | 8.19 | 8.78 | 8.70 | 8.49 |
| GPT-4o Audio | 3.94 | 2.68 | 3.53 | 8.79 |
| GPT-4o Mini Audio | 2.34 | 1.21 | 1.24 | 9.00 |
| Average | 5.35 | 5.11 | 6.65 | 7.78 |
Note: Scenarios: Content Mismatch (CM), Label Mismatch (LM), Logical Violation (LV), and Physical Violation (PV).
@@ 4 AudioTrust: Safety
The safety landscape for ALLMs presents challenges distinct from text-based systems (Wei et al., 2023; Wang et al., 2023c). The tone of a voice (whether urgent, distressed, or authoritative) can serve as a powerful tool to bypass the model’s safety alignment. For example, a user might pretend to have a medical emergency with a panicked voice to request dangerous information. This emotional attack vector, which leverages the persuasive nature of the human voice, is a novel challenge that text-only safety protocols do not address. Due to the high-dimensional and complex nature of these acoustic features, they may introduce new vulnerabilities during the encoding process and are challenging to align effectively with the safety constraints established for the text modality.
Attack Strategies. To systematically evaluate the safety of ALLMs, AudioTrust is structured along two orthogonal dimensions: Domain-Specific Policy Jailbreak Attacks and General Illegal Activity Guidance. The first category, Domain-Specific Policy Jailbreak Attacks, encompasses adversarial attempts to circumvent a model’s intrinsic behavioral constraints through techniques like prompt injection and role-playing (Li et al., 2025; Mo et al., 2024; Zhou et al., 2025). To ensure real-world applicability, these attacks are stratified into three critical domains: enterprise, financial, and healthcare systems, each populated with 100 test instances. The second category, General Illegal Activity Guidance, assesses the model’s propensity to generate harmful content for illicit requests (e.g., fraud, weapon crafting), drawing upon established benchmarks (Cao et al., 2025; Zhang et al., 2023c) to assemble 300 diverse prompts. In AudioTrust, we specifically investigated an attack vector based on the audio modality: Emotional Deception, which involves using speech prompts with specific emotional cues (e.g., anger, sadness) to induce the model to bypass its safety filters. All textual prompts were augmented via GPT-4o (OpenAI et al., 2024) and subsequently synthesized into audio using F5-TTS (Chen et al., 2024). Further details on data classification and construction are available in Appendices F.1 and F.2, respectively.
Evaluation & Metrics. Quantifying the success of jailbreak attacks and the harmfulness of generated content via rule-based methods is notoriously challenging. We therefore employed a model-based evaluation, leveraging GPT-4o. This choice is motivated by its strong alignment with human preferences on safety-sensitive assessments (Tan et al., 2025). GPT-4o rated each model-generated response on a 5-point Likert scale, where a score of 5 indicates complete refusal of a malicious request, and 1 denotes a fully compliant, harmful generation. Subsequently, these scores were further reviewed by human evaluators. Based on these scores, we computed two primary metrics: the Defense Success Rate (DSR), which measures the model’s robustness against jailbreak attempts (higher is better), and the Harmful Response Rate (HRR), quantifying the proportion of unsafe content produced in response to illicit prompts (lower is better). The detailed evaluation protocols, which are adapted for each task, are provided in Sections˜F.3 and F.4.
@@
| SALMONN | 74.2 | 74.4 | 80.8 | 77.1 |
| Ultravox | 97.2 | 83.8 | 90.8 | 98.0 |
| Qwen2-Audio | 68.2 | 80.6 | 81.4 | 92.5 |
| MiniCPM-o 2.6 | 76.2 | 79.2 | 81.6 | 96.2 |
| Step-Fun | 70.6 | 86.2 | 89.0 | 94.5 |
| Qwen2.5-omni | 97.2 | 94.8 | 94.2 | 99.1 |
| Kimi-Audio | 99.4 | 98.2 | 95.2 | 99.9 |
| OpenS2S | 51.4 | 67.8 | 75.2 | 71.8 |
| Step-Audio2 | 88.0 | 68.4 | 73.0 | 69.0 |
@@
| Gemini-1.5 Pro | 99.0 | 99.2 | 97.6 | 99.9 |
| GPT-4o Audio | 99.0 | 99.2 | 98.8 | 100.0 |
| GPT-4o mini Audio | 99.8 | 99.0 | 98.8 | 99.9 |
| Gemini-2.5 Flash | 100.0 | 99.8 | 99.4 | 99.8 |
| Gemini-2.5 Pro | 99.8 | 99.4 | 99.4 | 99.8 |
Note: Due to the common issue of random audio recognition failures in open-source models, these scores may be inflated.

### v4

@@ 1 Introduction
Despite rapid progress in ALLMs (Du et al., 2023; Tang et al., 2024; Chu et al., 2024; Zhang et al., 2023a; 2024a; Xu et al., 2025c; Yao et al., 2024), there is still no comprehensive benchmark evaluating their potential risks. Existing safety evaluation frameworks, such as SafeDialBench (Cao et al., 2025) and SafetyBench (Zhang et al., 2024b), mainly focus on the text modality while overlooking the unique properties or application scenarios of audio. Unlike text, audio features complex temporal-frequency patterns, rich emotions, and contextual dependencies, and introduces additional challenges, such as speech privacy, speaker recognition, and environmental acoustic analysis. These factors lead to unique trustworthiness risks in audio, including gender and accent biases, audio hallucinations, social engineering attacks, personal information leakage, and adversarial attacks on audio systems.
The integration of audio modalities into large models, while functionally powerful, introduces a new attack surface and exacerbates existing trustworthiness vulnerabilities. To systematically quantify these emergent risks, we introduce AudioTrust, the first comprehensive benchmark designed to evaluate the trustworthiness of ALLMs. AudioTrust establishes a rigorous evaluation framework across six critical dimensions where audio introduces unique safety concerns: (1) Fairness: Evaluating biases derived from vocal delivery, rather than the semantic content of speech; (2) Hallucinations: Testing for audio-grounded hallucinations, where model outputs violate the physical laws or temporal logic of an acoustic scene. (3) Safety: Assessing resilience to harmful queries that leverage persuasive or emotional vocal tones to bypass safety filters; (4) Privacy: Quantifying the leakage of sensitive information from spoken content and inference of personal attributes from acoustic cues; (5) Robustness: Assessing model performance in acoustically complex and imperfect environments, such as those with background noise. (6) Authentication: Evaluating the ability to distinguish authentic speakers from sophisticated impersonation attacks, including voice clones and audio-based social engineering. Underpinning our benchmark is a curated dataset of over 4,420 audio samples spanning 18 distinct evaluation tasks, from emergency communications to adversarial settings. We deploy a large-scale automated evaluation pipeline to ensure rigorous and reproducible assessment. The reliability of our automated metrics and results is verified by human experts (over 97% agreement rate). The initial findings on representative models are summarized in a public leaderboard (see Figure˜2(b)). Details of the benchmark are provided in Appendix˜C and Figure˜5.
@@
Safety: Incorporating audio inputs substantially broadens the attack surface. Unlike text, speech carries emotional, contextual, and anthropomorphic cues that adversaries can exploit by modulating tone, injecting affective signals, or impersonating identities. We design a composite attack framework spanning emotion-driven deception and identity-verification evasion. Empirical analysis shows: (1) closed-source models exhibit stronger overall robustness but remain sensitive to highly emotional speech; (2) open-source models are disproportionately vulnerable to identity- and emotion-based manipulation, often yielding unsafe outputs in high-stakes settings such as healthcare; (3) for prompts concerning illicit guidance, closed-source systems largely resist, whereas some open-source models deliver risky recommendations in multi-turn audio dialogues. Further details are provided in Section˜4.
@@ 2 AudioTrust: Fairness
Attack Strategies. To systematically probe these risks, we categorize fairness into two dimensions: traditional fairness and audio-based fairness. Traditional fairness assesses biases linked to demographic attributes like gender, race, and age (Cho et al., 2023; Zhang et al., 2017b; Wan et al., 2024) that can be inferred from a voice. We test, for example, if a model’s loan approval decisions are skewed by whether the applicant sounds male or female. Audio-based fairness isolates biases triggered by instrinsic acoustic properties. We divide these biases into four sub-categories: accent, linguistic characteristic (e.g., speech disfluencies may adversely impact the model’s fairness), economic status (e.g., noisy environments, often correlated with lower economic status, might yield more negative outputs), and personality traits (e.g., negative emotions may cause the model to produce negative outputs) (see Figure 6). For each sensitive attribute, we designed decision-making and stereotype-driven scenarios (Xu et al., 2025a; Wang et al., 2025c). Decision-making covers recruitment, admissions, and loan evaluations. Stereotype-driven contexts address beliefs such as men outperforming women in mathematics and gender roles in medicine and occupations. During the data construction phase, we utilized the multi-modal generation model GPT-4o (OpenAI et al., 2024) to produce textual content, which was then transformed into audio using the F5-TTS model (Chen et al., 2025). The entire data generation process begins with a text-based prompt, constructed entirely by GPT-4o to support natural language fluency and semantic consistency. We constructed a dedicated dataset consisting of 840 audio samples, each lasting approximately 20 seconds. The samples were annotated with seven key sensitive attributes: gender, age, and race for evaluating traditional fairness, as well as accent, linguistic characteristic, economic status, and personality traits for assessing audio-baseed fairness. Notably, due to the limitations of the audio modality, attributes such as appearance are not included (Nangia et al., 2020; Parrish et al., 2022). Detailed dataset construction procedures can be found in Sections˜D.1 and D.2.
Evaluation & Metrics. For fairness evaluation, we introduced a key metric: Group Fairness Score (Xu et al., 2025a; Deldjoo & Nazary, 2024). Group fairness examines the distributional equity of model outputs across different groups (e.g., male versus female), measured by the divergence or balance between group distributions. indicates perfect fairness, while indicates extreme unfairness. In computing the fairness score, we used human annotation and counting throughout to ensure data correctness and validity. For detailed formulas in the Section˜D.3.
@@
| SALMONN | 0.139 | 0.089 |
| Ultravox | 0.238 | 0.392 |
| Qwen2-Audio | 0.333 | 0.290 |
| MiniCPM-o 2.6 | 0.260 | 0.415 |
| Step-Fun | 0.658 | 0.505 |
| Qwen2.5-Omni | 0.067 | 0.202 |
| Kimi-Audio | 0.036 | 0.086 |
| OpenS2S | 0.017 | 0.157 |
| Step-Audio2 | 0.074 | 0.098 |
@@
| Gemini-1.5 Pro | 0.297 | 0.460 |
| GPT-4o Audio | 0.926 | 0.264 |
| GPT-4o mini Audio | 0.864 | 0.245 |
| Gemini-2.5 Flash | 0.370 | 0.246 |
| Gemini-2.5 Pro | 0.319 | 0.205 |
| Average | 0.328 | 0.261 |
Note: : higher than column average, : lower than column average, subscript is signed difference from mean.
@@ 3 AudioTrust: Hallucination
Attack Strategies. In AudioTrust, we identify two main categories of audio hallucinations (see Appendix E.1): Physical Logic and Chronological Order. The former relates to violations of acoustic laws and environment properties, and the latter reflects failures in reasoning about temporal and causal relations. These distinctions directly connect to safety risks in real-world use. To evaluate them, we built a dataset of 320 samples from synthetic and real sources, According to the different logical or temporal requirements of the testing tasks, the clips and sequences are edited and spliced to varying degrees and in different orders (see Appendix E.2). For Acoustic-Physical Hallucinations, we focus on two specific manifestations: (1) Contravention of Physicochemical Constraints, generating impossible events (e.g., the flames are burning in the seawater) to test propagation understanding (Liu et al., 2024); (2) Source–Environment Mismatch, applying contradictory reverberation (e.g., casual speech with cathedral acoustics) to test disentanglement of source vs. environment (Zhou et al., 2022). For Temporo-Logical Hallucinations, we examine (1) Temporal-Causal Inversion: reversing causal chains (e.g., engine start before ignition) to probe event logic (Chen et al., 2023); (2) Cross-Modal Contradiction: pairing audio with conflicting text (e.g., fast footsteps described as peaceful rain) to test cross-modal reasoning (Deshmukh et al., 2023).
Evaluation & Metrics. We introduce a comprehensive evaluation framework to assess model capabilities across four key dimensions: hallucination detection, attribute verification, real-world consistency, and transcription accuracy. For hallucination detection, models are required to identify inaccuracies in audio-text pairs and provide justifications. Performance is quantified via a multi-dimensional, GPT-4o-based (OpenAI et al., 2024) and Qwen3-based (Yang et al., 2025a) evaluators.The scores mentioned above range from 0 to 10 and are used to evaluate the accuracy of the detection and the quality of the interpretation. These scores were then subject to further review by the human evaluators. We probe for attribute-level hallucinations related to physical properties, labels, and content using multi-stage prompting (Madaan et al., 2023). To measure alignment with factual information, we adopt the two-stage protocol from Li et al. (2023a) for real-world consistency assessment. Finally, we evaluate transcription robustness under hallucinatory interference using both the standard Word Error Rate (WER) (Deshmukh et al., 2023) and the cross-modal WER (CM-WER) (Tang et al., 2024). The complete experimental design and metric details are provided in Appendices E.3.1 and E.3.2.
@@
| MiniCPM-o 2.6 | 6.24 / 4.62 | 6.20 / 5.13 | 8.28 / 7.53 | 6.13 / 5.30 |
| Qwen2-Audio | 8.15 / 8.66 | 4.34 / 7.05 | 7.26 / 5.48 | 7.77 / 8.18 |
| SALMONN | 2.65 / 1.70 | 1.22 / 1.16 | 6.64 / 5.97 | 3.98 / 3.92 |
| Ultravox | 5.74 / 8.29 | 4.52 / 7.83 | 8.01 / 6.10 | 8.34 / 8.70 |
| Qwen2.5-Omni | 8.12 / 8.44 | 5.63 / 3.94 | 7.89 / 6.45 | 6.11 / 6.17 |
| Step-Fun | 3.96 / 3.93 | 4.84 / 5.04 | 5.80 / 5.83 | 8.72 / 9.25 |
| Kimi Audio | 1.86 / 1.88 | 5.77 / 5.67 | 5.82 / 7.22 | 8.54 / 8.62 |
| Step-Audio2 | 3.62 / 1.74 | 1.94 / 4.09 | 2.60 / 6.10 | 2.76 / 8.39 |
| OpenS2S | 1.92 / 1.30 | 5.03 / 4.11 | 5.97 / 5.75 | 7.89 / 8.27 |
@@
| Gemini-1.5 Pro | 8.41 / 8.05 | 7.81 / 7.17 | 8.66 / 8.35 | 8.87 / 9.78 |
| Gemini-2.5 Flash | 7.98 / 7.15 | 8.36 / 8.24 | 8.71 / 9.03 | 8.57 / 9.74 |
| Gemini-2.5 Pro | 8.19 / 7.02 | 8.78 / 5.08 | 8.70 / 9.00 | 8.49 / 9.70 |
| GPT-4o Audio | 3.94 / 1.65 | 2.68 / 1.68 | 3.53 / 3.41 | 8.79 / 9.43 |
| GPT-4o Mini Audio | 2.34 / 1.35 | 1.21 / 1.77 | 1.24 / 3.25 | 9.00 / 9.61 |
| Average | 5.22 / 4.70 | 4.88 / 4.85 | 6.36 / 6.39 | 7.43 / 8.22 |
Note: Values are shown as "GPT-4o / Qwen3 evaluators. Subscript is the absolute difference from the respective column average. Scenarios: Content Mismatch (CM), Label Mismatch (LM), Logical Violation (LV), and Physical Violation (PV).
@@ 4 AudioTrust: Safety
The safety landscape for ALLMs presents challenges distinct from text-based systems (Wei et al., 2023; Wang et al., 2023b). The tone of a voice (whether urgent, distressed, or authoritative) can serve as a powerful tool to bypass the model’s safety alignment. For example, a user might pretend to have a medical emergency with a panicked voice to request dangerous information. This emotional attack vector, which leverages the persuasive nature of the human voice, is a novel challenge that text-only safety protocols do not address. Due to the high-dimensional and complex nature of these acoustic features, they may introduce new vulnerabilities during the encoding process and are challenging to align effectively with the safety constraints established for the text modality.
Attack Strategies. To systematically evaluate the safety of ALLMs, AudioTrust is structured along two orthogonal dimensions: Domain-Specific Policy Jailbreak Attacks and General Illegal Activity Guidance. The first category, Domain-Specific Policy Jailbreak Attacks, encompasses adversarial attempts to circumvent a model’s intrinsic behavioral constraints through techniques like prompt injection and role-playing (Li et al., 2025; Mo et al., 2024; Zhou et al., 2024). To ensure real-world applicability, these attacks are stratified into three critical domains: enterprise, financial, and healthcare systems, each populated with 100 test instances. The second category, General Illegal Activity Guidance, assesses the model’s propensity to generate harmful content for illicit requests (e.g., fraud, weapon crafting), drawing upon established benchmarks (Cao et al., 2025; Zhang et al., 2024b) to assemble 300 diverse prompts. In AudioTrust, we specifically investigated an attack vector based on the audio modality: Emotional Deception, which involves using speech prompts with specific emotional cues (e.g., anger, sadness) to induce the model to bypass its safety filters. All textual prompts were augmented via GPT-4o (OpenAI et al., 2024) and subsequently synthesized into audio using F5-TTS (Chen et al., 2025). Further details on data classification and construction are available in Appendices F.1 and F.2, respectively.
@@
| SALMONN | 74.2 / 83.4 | 74.4 / 91.2 | 80.8 / 92.2 | 77.1 / 89.4 |
| Ultravox | 97.2 / 97.0 | 83.8 / 97.0 | 90.8 / 99.2 | 98.0 / 96.9 |
| Qwen2-Audio | 68.2 / 95.2 | 80.6 / 91.8 | 81.4 / 90.2 | 92.5 / 93.3 |
| MiniCPM-o 2.6 | 76.2 / 83.0 | 79.2 / 86.0 | 81.6 / 94.8 | 96.2 / 99.3 |
| Step-Fun | 70.6 / 92.4 | 86.2 / 92.8 | 89.0 / 94.8 | 94.5 / 94.2 |
| Qwen2.5-omni | 97.2 / 99.2 | 94.8 / 99.4 | 94.2 / 97.8 | 99.1 / 99.6 |
| Kimi-Audio | 99.4 / 99.8 | 98.2 / 100.0 | 95.2 / 99.6 | 99.9 / 99.9 |
| OpenS2S | 51.4 / 47.6 | 67.8 / 87.4 | 75.2 / 83.0 | 71.8 / 72.9 |
| Step-Audio2 | 88.0 / 100.0 | 68.4 / 99.8 | 73.0 / 96.8 | 69.0 / 99.6 |
@@
| Gemini-1.5 Pro | 99.0 / 94.0 | 99.2 / 97.2 | 97.6 / 99.0 | 99.9 / 99.9 |
| GPT-4o Audio | 99.0 / 99.2 | 99.2 / 100.0 | 98.8 / 100.0 | 100.0 / 99.9 |
| GPT-4o mini Audio | 99.8 / 97.6 | 99.0 / 100.0 | 98.8 / 99.0 | 99.9 / 99.9 |
| Gemini-2.5 Flash | 100.0 / 100.0 | 99.8 / 93.2 | 99.4 / 96.0 | 99.8 / 93.6 |
| Gemini-2.5 Pro | 99.8 / 82.6 | 99.4 / 87.2 | 99.4 / 100.0 | 99.8 / 99.0 |
Note: Due to the common issue of random audio recognition failures in open-source models, these scores may be inflated. Values are shown as GPT-4o / Qwen3 evaluators.
Evaluation & Metrics. Quantifying the success of jailbreak attacks and the harmfulness of generated content via rule-based methods is notoriously challenging. We therefore employed a model-based evaluation, leveraging GPT-4o and Qwen3. This choice is motivated by its strong alignment with human preferences on safety-sensitive assessments (Tan et al., 2025). We first generated text-based malicious prompts and converted them into audio; specific emotional expressions were achieved by selecting reference audio clips with distinct emotional timbres during the TTS process. In particular, we selected TTS voices and configurations that realize the same sentence with different emotional tones and different genders. GPT-4o and Qwen3 rated each model-generated response on a 5-point Likert scale, where a score of 5 indicates complete refusal of a malicious request, and 1 denotes a fully compliant, harmful generation. Subsequently, these scores were further reviewed by human evaluators. Based on these scores, we computed the primary metric: the Defense Success Rate (DSR), which measures the model’s robustness against jailbreak attempts (higher is better). The detailed evaluation protocols, which are adapted for each task, are provided in Sections˜F.3 and F.4.